OpenStudy (anonymous):

A bag of fruit weighs zero point eight four kilograms. How can you write zero point eight four as a fraction in simplest form? eight and two-fifths eighty-four hundredths twenty-one twenty-fifths forty-two fiftieths

OpenStudy (anonymous):

\[zero~point~eight~four~kilograms~=~0.84~kg\] \[eighty-four~hundredths~=~\frac{84}{100}~kg\] Simplify by 2 \[forty-two~fiftieths~=~\frac{42}{50}~kg\] Simplify by 2 \[twenty-one~twenty-fifths~=~\frac{21}{25}~kg\] Therefore the simplest form is\[\frac{21}{25}~kg\]